OpenStudy (anonymous):

A farmer has 268 cows and each cow has either 1 or 2 calves. If there are 335 calves, what fraction of the cows have 2 calves?

OpenStudy (agent0smith):

I'll use S for a cow that has a single calf, and T for a cow that has two calves. There's 268 cows total, so : S+T = 268 There's 335 calves total, so 1*S + 2*T will be equal to 335. S+2T = 335

OpenStudy (agent0smith):

So now you have a system of two equations and two variables to solve: S+T = 268 S+2T = 335 Try solving this for T, since we want the fraction with 2 calves. And remember the fraction means the number of cows with two calves over the total number of cows, or T/(S+T)

OpenStudy (agent0smith):

ie T/268 will be your fraction

OpenStudy (anonymous):

Thanks so much for clearing this up for me. And I got my answer... If each cow has one calf, then there are a total of 268 calves. So we take the given number and minus it from the total number of calves. 335-268=67. 67 cows have 2 calves. Our fraction is 67/268, which, when simplifies, is; 1/4.

OpenStudy (agent0smith):

Yep :) Your method is actually a bit easier than the way I showed you.